OpenStudy (anonymous):

Use Intermediate Value Theorem to prove that there exists some number c on [0,2] such that f(c)=0. g(x)= 3^x-4/x

OpenStudy (anonymous):

i would first evaluate at the endpoints g(0) and g(2), if they are 0 then you have your answer, if they arent 0 but one is < 0 and one is > 0 and g(x) is continuous then there exists a number c such that f(c) = 0

OpenStudy (anonymous):

You can't put zero under the 4 though

OpenStudy (anonymous):

i meant if g(x) is continuous on the interval [0,2]

OpenStudy (anonymous):

I tried putting .01 instead of 0

OpenStudy (anonymous):

oh, lol, my bad

OpenStudy (anonymous):

forgot about divide by 0

OpenStudy (anonymous):

Can you do .01 instead of 0?

OpenStudy (anonymous):

since it is still a number b/w the interval?

OpenStudy (anonymous):

is this for calc 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

have you learned local min an max?

OpenStudy (anonymous):

not that i recall

OpenStudy (anonymous):

ok you dont need to test the endpoints, if you can prove it within a smaller interval that is in the interval [0,2] then it is sufficient

OpenStudy (anonymous):

so prove it for the interval [1,2]

OpenStudy (anonymous):

and I think it should be f(x)= 3^x-4/x not g(x)= 3^x-4/x, right?